# On the geometry of the free factor graph for ${\rm{Aut}}(F_{N})$

### 06 Dec 2023 v1

@@ Abstract
Let $\Phi$ be a pseudo-Anosov diffeomorphism of a compact (possibily non-orientable) surface $\Sigma$ with one boundary component. We show that if $b\in\pi_{1}(\Sigma)$ is the boundary word, $\phi\in{\rm{Aut}}(\pi_{1}(\Sigma))$ is a representative of $\Phi$ fixing $b$, and ${\rm{ad}}_{b}$ denotes conjugation by $b$, then the orbits of $\langle\phi,{\rm{ad}}_{b}\rangle\cong\mathbb{Z}^{2}$ in the graph of free factors of $\pi_{1}(\Sigma)$ are quasi-isometrically embedded. It follows that for $N\geq 2$ the free factor graph for ${\rm{Aut}}(F_{N})$ is not hyperbolic, in contrast to the ${\rm{Out}}(F_{N})$ case.
@@ 1 Introduction
The natural map ${\rm{Out}}(F_{N})\to{\rm{Isom}}(\mathcal{OF}_{N})$ is an isomorphism [BB23], so Theorem A describes all possible actions of cyclic groups on $\mathcal{OF}_{N}$. As the graph is hyperbolic, we also know that $\mathbb{Z}^{2}$ cannot with act quasi-isometric orbits. In fact, results in the literature [BB23, HM20] tell us something much stronger:
@@
Let $\Sigma$ be a compact surface with one boundary component, fix an identification $\pi_{1}\Sigma=F_{N}$, let $b\in F_{N}$ be the homotopy class of the boundary loop, let $\phi\in{\rm{Aut}}(F_{N})$ be the automorphism induced by a pseudo-Anosov element of $\text{{\rm{Mod}}}(\Sigma)$, let $\Lambda=\langle\phi,\,{\rm{ad}}_{b}\rangle<{\rm{Aut}}(F_{N})$ and note that $\Lambda\cong\mathbb{Z}^{2}$. Then, for every $A\in\mathcal{AF}_{N}$, the orbit map $A\mapsto\lambda(A)$, defines a quasi-isometric embedding $\Lambda\cong\mathbb{Z}^{2}\hookrightarrow\mathcal{AF}$.
@@
A crucial property of the boundary element $b\in F_{N}$ is that it is filling in the sense that it is not contained in any proper free factor of $F_{N}$. Given any filling element $b$, we analyse the action of ${\rm{ad}}_{b}$ on $\mathcal{AF}_{N}$ with the aim or showing that the orbits of $\langle{\rm{ad}}_{b}\rangle$ give quasi-isometric embeddings of $\mathbb{Z}$. The key tool in this analysis is the notion of $b$-reduced decomposition for elements $w\in F_{N}$, which we introduce in Section 3. Using these decompositions, we define an integer-valued invariant $[A]_{b}$ of free factors $A$ and we use this to define a Lipschitz retraction of $\mathcal{AF}_{N}$ onto the $\langle{\rm{ad}}_{b}\rangle$-orbit of a vertex $V$; roughly speaking, this retraction sends the vertex $A$ to ${\rm{ad}}_{b}^{n}(V)$ if $n=[A]_{b}$.
@@
Acknowledgements. This paper is dedicated to Slava Grigorchuk on his 70th birthday, with thanks for his many highly-original contributions to group theory. The first author acknowledges the support of the National Science Foundation under the grant number DMS-2304774. The third author is supported by the Royal Society of Great Britain through a University Research Fellowship.
@@ 2.1 Whitehead graphs, word length, and the cut-vertex lemma
An element of $F_{N}$ is primitive if it belongs to some basis, and is simple if it belongs to some proper free factor. Every element $w$ has a cyclic reduction $w_{0}$, obtained by writing $w\equiv w_{1}w_{0}w_{1}^{-1}$ with $w_{1}$ as long as possible. An element is cyclically reduced (with respect to the basis $\mathcal{S}$) if $w=w_{0}$. We use ${\mathrm{Wh}}_{\mathcal{S}}(w)$ to denote the cyclic Whitehead graph of $w$. This is the graph with vertex set $\mathcal{S}\cup\mathcal{S}^{-1}$ that has an edge from $x$ to $y^{-1}$ if $xy$ is a subword of the cyclic reduction $w_{0}$ or if $x$ is the last letter of $w_{0}$ and $y$ is the first letter of $w_{0}$. A vertex $v$ of a graph is a cut vertex if the full subgraph spanned by vertices not equal to $v$ is disconnected. The famous lemma below was stated by Whitehead for primitive elements [Whi36] and the generalization to simple elements was observed by various authors (see, e.g. [Mar95, Proposition 49] or [Ota92, Sta99, HW19]). (For readers who pursue the references: Whitehead uses the word simple differently to modern authors—his simple sets are subsets of bases.)
@@ 3.1 $b$-reduced decompositions and a cancellation lemma
Fix a basis $\mathcal{S}$ of $F_{N}$. Given a cyclically reduced word $b$ and a reduced word $w$, representing elements of $F_{N}$, the $b$-reduced decomposition of $w$ (with respect to $\mathcal{S}$) is the decomposition
@@
Let $T$ be the Cayley tree for $F_{N}$ with respect to the basis $\mathcal{S}$. As $b$ is cyclically reduced, the axis $X_{b}$ of $b$ passes through the identity vertex $1\in T$. Let $X_{a}$ be the axis for $a$. Let $p$ be the shortest path from $1$ to $X_{a}$. The intersection of $p$ with $X_{b}$ is of length at most $|b|_{\mathcal{S}}-1$, as the reduced word representing $a$ starts with the word labelling $p$ and ends with its inverse and $[a]_{b}=0$. Furthermore, the intersection of $X_{a}$ with $X_{b}$ can have length at most $|b|_{\mathcal{S}}$: to see this, note that the Whitehead graph ${\mathrm{Wh}}_{\mathcal{S}}(b)$ does not contain a cut vertex by Proposition 2.2, whereas ${\mathrm{Wh}}_{\mathcal{S}}(a)$ does by the Cut Vertex Lemma, so $X_{a}$ cannot contain all the turns (pairs of consecutive of labels from $\mathcal{S}\cup\mathcal{S}^{-1}$) that appear along $X_{b}$.
@@
Suppose $b\in F_{N}$ is cyclically reduced with respect to ${\mathcal{S}}$ and that the Whitehead graph ${\mathrm{Wh}}_{\mathcal{S}}(b)$ is connected with no cut vertex. Let $T$ be the Cayley tree with respect to $\mathcal{S}$ and let $X_{b}$ be the axis of $b$ in $T$. If $A\leq F_{N}$ is a proper free factor and $T_{A}$ is the minimal subtree of $A$ in $T$, then
@@
Proposition 2.2 tells us that Proposition 3.5 applies to $b$ and Proposition 3.5(2) assures us that if $A<B$ then $|[A]^{\mathcal{S}}_{b}-[B]^{\mathcal{S}}_{b}|\leq 1$; this covers the case $N=3$. For the case $N=2$, a different argument is needed, because $\mathcal{AF}_{N}$ is defined differently. In this case, the vertex $\langle a_{1}\rangle$ is adjacent to $\langle a_{2}\rangle$ if $\{a_{1},a_{2}\}$ is a basis. As in Proposition 3.5(1), we know that the projection $\pi_{X_{b}}(X_{a_{i}})$ of each axis to $X_{b}$ has length at most $|b|_{\mathcal{S}}$. We claim that furthermore, if the projections are disjoint then the arc connecting $\pi_{X_{b}}(X_{a_{1}})$ to $\pi_{X_{b}}(X_{a_{2}})$ is length at most $|b|_{\mathcal{S}}$. This claim clearly gives a uniform bound on the difference between $[a_{1}]_{b}$ and $[a_{2}]_{b}$, and with a similar case-by-case analysis as in Proposition 3.5 one can check that $[a_{1}]_{b}$ and $[a_{2}]_{b}$ differ by at most $2$.
@@ 3.2 Change of Basis
Let $T^{\mathcal{S}}$ and $T^{\mathcal{T}}$ be the Cayley trees of $F_{N}$ given by $\mathcal{S}$ and $\mathcal{T}$, respectively. Let $f:T^{\mathcal{S}}\to T^{\mathcal{T}}$ be the map that sends each edge of $T^{\mathcal{S}}$ to the geodesic in $T^{\mathcal{T}}$ with the same endpoints. If $p=[1,g]_{\mathcal{S}}$ is a path in $T^{\mathcal{S}}$ from 1 to the axis of an element $a\in F_{N}$, then the bounded backtracking property (see [Coo87, GJLL98]) implies that $g$ lies at distance at most $C$ from the axis of $a$ in $T^{\mathcal{T}}$, where $C$ is the bounded backtracking constant. Therefore, as $|b|_{\mathcal{S}}=|b|_{\mathcal{T}}$, the proposition reduces to showing that for any path $p$ based at 1, the reduction of $f(p)$ spends approximately the same amount of time travelling along the axis of $b$ in $T^{\mathcal{T}}$ as the path $p$ does along the axis of $b$ in $T^{\mathcal{S}}$ (in the same direction). This is reasonably standard: decompose such a path $p$ as $p=p_{1}p_{2}p_{3}$, where $p_{1}=[1,b^{k}]_{\mathcal{S}}$ for some $k$, the path $p_{2}$ is a segment of length $\leq|b|_{\mathcal{S}}$, and $p_{3}$ is a path disjoint from the axis $X^{\mathcal{S}}_{b}\subset T^{\mathcal{S}}$ except for its initial endpoint. As $f$ is bi-Lipschitz on the vertex set, $p_{3}$ can further be decomposed as $p_{3}^{\prime}p_{3}^{\prime\prime}$, where $f(p_{3}^{\prime\prime})$ is disjoint from $X^{\mathcal{T}}_{b}$ and $p_{3}^{\prime}$ is of uniformly bounded length. Then the intersection of the reduction of $f(p)$ with $X^{\mathcal{T}}_{b}$ in $T_{\mathcal{T}}$ is $[1,b^{k}]_{\mathcal{T}}$, up to cancellation/concatenation with a subpath of $f(p_{2}p_{3}^{\prime})$, which is of uniformly bounded length. ∎
@@ 4 Building quasi-flats
In Sections 1-3 of [BH92], it is shown that an automorphism $\phi_{0}$ as above has a stable train track representative $f\colon G\to G$ with exactly one indivisible Nielsen path (iNP) $\rho$ which, as the automorphism is geometric, forms a loop representing $b$ based at a point $x\in G$ that is fixed by $f$. Furthermore, $\rho$ crosses every edge in $G$ exactly twice (see [BH92, Lemma 3.9] and the discussion in [BH92, Section 4]). Identifying $\pi_{1}(G,x)$ with $F_{N}$, we define $\phi$ to be the automorphism induced by $f_{*}\colon\pi_{1}(G,x)\to\pi_{1}(G,x)$ and take $\mathcal{S}$ to be a basis of $F_{N}$ determined by a choice of maximal tree $T$ in $G$ together with an orientation of the edges in $G\smallsetminus T$. Each edge-loop based at $x$ determines a word in the letters ${\mathcal{S}}^{\pm 1}$ whose length is the number of edges of $G\smallsetminus T$ crossed by the loop. In particular, $|b|_{\mathcal{S}}=2N$. Since $b$ is the boundary for a surface of Euler characteristic $1-N$, this tells us that $\mathcal{S}$ minimizes the length of $b$.
We now prove the second part with $r\geq 0$. There exists a legal loop $l$ in $G$ based at $x$ that crosses some edge exactly once (this is a variant of Francaviglia and Martino’s Sausage Lemma [FM11, Lemma 3.14]). Any such loop represents a primitive element in $F_{N}$. As in Section 2 of [BH92], we lift $f$ to a map $\tilde{f}\colon\tilde{G}\to\tilde{G}$ on the universal cover that represents $\phi$ (i.e. $\tilde{f}(gy)=\phi(g)\tilde{f}(y)$ for all $y\in\tilde{G}$). The map $\tilde{f}$ fixes a lift $\tilde{x}$ of our basepoint $x\in G$, and the axis of $b$ (acting as a deck transformation of $\tilde{G}$) contains $\tilde{x}$. As the loop $l$ is based at $x$ its axis $\operatorname{Axis}_{\tilde{G}}(l)$ in $\tilde{G}$ also contains $\tilde{x}$. As $l$ is legal and $\tilde{x}$ is fixed by $\tilde{f}$, the axis of $\phi^{r}(l)$ also contains $\tilde{x}$ for all $r\geq 0$. It follows that iterated images of $l$ are cyclically reduced, so that $[\phi^{r}(l)]_{b}=0$ for all $r\geq 0$. For $k\in\mathbb{Z}$, the axis of ${\rm{ad}}_{b}^{k}(l)$ is also legal and contains the point $b^{k}\tilde{x}$, so that $k-1\leq[{\rm{ad}}_{b}^{k}(l)]_{b}\leq k+1$. As $b^{k}\tilde{x}$ is also fixed under $\tilde{f}$, this implies that $k-1\leq[\phi^{r}{\rm{ad}}_{b}^{k}(l)]_{b}\leq k+1$ for all $k\in\mathbb{Z}$ and $r\geq 0$. As the map $\mathcal{AF}_{N}\to\mathbb{Z}$ induced by $[A]\mapsto[A]_{b}$ is 2-Lipschitz, this implies that for all $A\in\mathcal{AF}_{N}$ there exists $C_{A}$ such that $|[\phi^{r}{\rm{ad}}_{b}^{k}(A)]_{b}-[{\rm{ad}}_{b}^{k}(A)]_{b}|\leq C_{A}$ for all $r\geq 0$ and $k\in\mathbb{Z}$.
@@
To deal with the case where $r<0$ in (2), we run the entire argument again for $[\phi_{0}]^{-1}$. This provides a representative $\psi$ of $[\phi_{0}]^{-1}$ fixing $b$, such that for all $A\in\mathcal{AF}_{N}$ there exists a constant $C_{A}^{\prime}$ such that $|[\psi^{r}{\rm{ad}}_{b}^{k}(A)]_{b}-[{\rm{ad}}_{b}^{k}(A)]_{b}|\leq C_{A}^{\prime}$ for all $k\in\mathbb{Z}$ and $r\geq 0$. Note that although you might be working with a different basis $\mathcal{S}^{\prime}$ on which $b$ is minimal length, the projections $\mathcal{AF}_{N}\to\mathbb{Z}$ are coarsely equivalent by Proposition 3.7. Furthermore, this is the unique representative of $[\phi^{\prime}]^{-1}$ fixing $b$ such that $b^{\infty}$ and $b^{-\infty}$ are nonattracting under $\partial\psi$. This implies that $\psi=\phi^{-1}$, so we can finish part (2) by taking $M_{A}:=\max\{C_{A},C_{A}^{\prime}\}$.
@@
Let $\Sigma$ be a compact surface with one boundary component, fix an identification $\pi_{1}\Sigma=F_{N}$, let $b\in F_{N}$ be the homotopy class of the boundary loop, let $\phi\in{\rm{Aut}}(F_{N})$ be the automorphism induced by a pseudo-Anosov element of $\text{{\rm{Mod}}}(\Sigma)$, let $\Lambda=\langle\phi,\,{\rm{ad}}_{b}\rangle<{\rm{Aut}}(F_{N})$ and note that $\Lambda\cong\mathbb{Z}^{2}$. Then, for every $A\in\mathcal{AF}_{N}$, the orbit map $A\mapsto\lambda(A)$, defines a quasi-isometric embedding $\Lambda\cong\mathbb{Z}^{2}\hookrightarrow\mathcal{AF}$.
@@
Note that it is enough to prove the result for one orbit. We may also replace $\phi$ with ${\rm{ad}}_{b}^{k}\phi$ without changing $\Lambda$, so we can assume that $\phi$ and the basis ${\mathcal{S}}$ are as described in Proposition 4.1. Let $s\in\mathcal{S}$ and let $A=\langle s\rangle$. Then $k-1\leq[{\rm{ad}}_{b}^{k}(A)]_{b}^{\mathcal{S}}\leq k+1$ for all $k\in\mathbb{Z}$. We define a map $\mathcal{OF}_{N}\to\mathbb{Z}$ as follows: for each conjugacy class $[B]$ of a free factor we pick a closest point $[\phi]^{j}([A])$ in the $[\phi]$-orbit of $[A]$. This assignment $[B]\mapsto j$ is Lipschitz [BF14], and by composing it with the natural projection $\mathcal{AF}_{N}\to\mathcal{OF}_{N}$ we obtain a Lipschitz map $R\colon\mathcal{AF}_{N}\to\mathbb{Z}$. It follows from Theorem 3.6 that the map $[\,\cdot\,]_{b}^{\mathcal{S}}\times R\colon\mathcal{AF}_{N}\to\mathbb{Z}^{2}$ is also Lipschitz, and we claim that $B\mapsto\phi^{R(B)}{\rm{ad}}_{b}^{[B]_{b}}(A)$ gives a Lipshitz retraction of $\mathcal{AF}_{N}$ onto the orbit of $\Lambda=\langle\phi,{\rm{ad}}_{b}\rangle$. To see this, note that $R(\phi^{j}{\rm{ad}}_{b}^{k}(A))=j$ for all $j,k\in\mathbb{Z}$, and

### 12 Mar 2024 v2

@@ Abstract
Let $\Phi$ be a pseudo-Anosov diffeomorphism of a compact (possibly non-orientable) surface $\Sigma$ with one boundary component. We show that if $b\in\pi_{1}(\Sigma)$ is the boundary word, $\phi\in{\rm{Aut}}(\pi_{1}(\Sigma))$ is a representative of $\Phi$ fixing $b$, and ${\rm{ad}}_{b}$ denotes conjugation by $b$, then the orbits of $\langle\phi,{\rm{ad}}_{b}\rangle\cong\mathbb{Z}^{2}$ in the graph of free factors of $\pi_{1}(\Sigma)$ are quasi-isometrically embedded. It follows that for $N\geq 2$ the free factor graph for ${\rm{Aut}}(F_{N})$ is not hyperbolic, in contrast to the ${\rm{Out}}(F_{N})$ case.
@@ 1 Introduction
The natural map ${\rm{Out}}(F_{N})\to{\rm{Isom}}(\mathcal{OF}_{N})$ is an isomorphism [BB23], so Theorem A describes all possible actions of cyclic groups on $\mathcal{OF}_{N}$. As the graph is hyperbolic, we also know that $\mathbb{Z}^{2}$ cannot act with quasi-isometric orbits. In fact, results in the literature [BB23, HM20] tell us something much stronger:
@@
Let $\Sigma$ be a compact surface with one boundary component, fix an identification $\pi_{1}\Sigma=F_{N}$, let $b\in F_{N}$ be the homotopy class of the boundary loop, let $\phi\in{\rm{Aut}}(F_{N})$ be the automorphism induced by a pseudo-Anosov element of $\text{{\rm{Mod}}}(\Sigma)$, let $\Lambda=\langle\phi,\,{\rm{ad}}_{b}\rangle<{\rm{Aut}}(F_{N})$ and note that $\Lambda\cong\mathbb{Z}^{2}$. Then, for every $A\in\mathcal{AF}_{N}$, the orbit map $\lambda\mapsto\lambda(A)$, defines a quasi-isometric embedding $\Lambda\cong\mathbb{Z}^{2}\hookrightarrow\mathcal{AF}$.
@@
A crucial property of the boundary element $b\in F_{N}$ is that it is filling in the sense that it is not contained in any proper free factor of $F_{N}$. Given any filling element $b$, we analyse the action of ${\rm{ad}}_{b}$ on $\mathcal{AF}_{N}$ with the aim of showing that the orbits of $\langle{\rm{ad}}_{b}\rangle$ give quasi-isometric embeddings of $\mathbb{Z}$. The key tool in this analysis is the notion of $b$-reduced decomposition for elements $w\in F_{N}$, which we introduce in Section 3. Using these decompositions, we define an integer-valued invariant $[A]_{b}$ of free factors $A$ and we use this to define a Lipschitz retraction of $\mathcal{AF}_{N}$ onto the $\langle{\rm{ad}}_{b}\rangle$-orbit of a vertex $V$; roughly speaking, this retraction sends the vertex $A$ to ${\rm{ad}}_{b}^{n}(V)$ if $n=[A]_{b}$.
@@
Acknowledgements. This paper is dedicated to Slava Grigorchuk on his 70th birthday, with thanks for his many highly-original contributions to group theory. We thank the anonymous referee for their helpful comments. The first author acknowledges the support of the National Science Foundation under the grant number DMS-2304774. The third author is supported by the Royal Society of Great Britain through a University Research Fellowship.
@@ 2.1 Whitehead graphs, word length, and the cut-vertex lemma
An element of $F_{N}$ is primitive if it belongs to some basis, and is simple if it belongs to some proper free factor. Every element $w$ has a cyclic reduction $w_{0}$, obtained by writing $w\equiv w_{1}w_{0}w_{1}^{-1}$ with $w_{1}$ as long as possible. An element is cyclically reduced (with respect to the basis $\mathcal{S}$) if $w=w_{0}$. We use ${\mathrm{Wh}}_{\mathcal{S}}(w)$ to denote the cyclic Whitehead graph of $w$. This is the graph with vertex set $\mathcal{S}\cup\mathcal{S}^{-1}$ that has an edge from $x$ to $y^{-1}$ if $xy$ is a subword of the cyclic reduction $w_{0}$ or if $x$ is the last letter of $w_{0}$ and $y$ is the first letter of $w_{0}$. Importantly for us, the cyclic Whitehead graph keeps track of the turns crossed by the axis of the element $w$ in the Cayley tree determined by $\mathcal{S}$.
A vertex $v$ of a graph is a cut vertex if the full subgraph spanned by vertices not equal to $v$ is disconnected. The famous lemma below was stated by Whitehead for primitive elements [Whi36] and the generalization to simple elements was observed by various authors (see, e.g. [Mar95, Proposition 49] or [Ota92, Sta99, HW19]). (For readers who pursue the references: Whitehead uses the word simple differently to modern authors—his simple sets are subsets of bases.)
@@ 3.1 $b$-reduced decompositions and a cancellation lemma
Fix a basis $\mathcal{S}$ of $F_{N}$. Given a cyclically reduced word $b$ and a reduced word $w$ representing elements of $F_{N}$, the $b$-reduced decomposition of $w$ (with respect to $\mathcal{S}$) is the decomposition
@@
As $\mathcal{S}$ minimizes $|b|_{\mathcal{S}}$, in particular $b$ is cyclically reduced. Let $T$ be the Cayley tree for $F_{N}$ with respect to the basis $\mathcal{S}$. As $b$ is cyclically reduced, the axis $X_{b}$ of $b$ passes through the identity vertex $1\in T$. Let $X_{a}$ be the axis for $a$. Let $p$ be the shortest path from $1$ to $X_{a}$. The intersection of $p$ with $X_{b}$ is of length at most $|b|_{\mathcal{S}}-1$, as the reduced word representing $a$ starts with the word labelling $p$ and ends with its inverse and $[a]_{b}=0$. Furthermore, the intersection of $X_{a}$ with $X_{b}$ can have length at most $|b|_{\mathcal{S}}$: to see this, note that the cyclic Whitehead graph ${\mathrm{Wh}}_{\mathcal{S}}(b)$ does not contain a cut vertex by Proposition 2.2, whereas ${\mathrm{Wh}}_{\mathcal{S}}(a)$ does by the Cut Vertex Lemma, so $X_{a}$ cannot contain all the turns (pairs of consecutive of labels from $\mathcal{S}\cup\mathcal{S}^{-1}$) that appear along $X_{b}$.
@@
Suppose $b\in F_{N}$ is cyclically reduced with respect to ${\mathcal{S}}$ and that the cyclic Whitehead graph ${\mathrm{Wh}}_{\mathcal{S}}(b)$ is connected with no cut vertex. Let $T$ be the Cayley tree with respect to $\mathcal{S}$ and let $X_{b}$ be the axis of $b$ in $T$. If $A\leq F_{N}$ is a proper free factor and $T_{A}$ is the minimal subtree of $A$ in $T$, then
@@
Proposition 2.2 tells us that Proposition 3.5 applies to $b$ and Proposition 3.5(2) assures us that if $A<B$ then $|[A]^{\mathcal{S}}_{b}-[B]^{\mathcal{S}}_{b}|\leq 1$; this covers the case $N=3$. For the case $N=2$, a different argument is needed, because $\mathcal{AF}_{N}$ is defined differently. In this case, the vertex $\langle a_{1}\rangle$ is adjacent to $\langle a_{2}\rangle$ if $\{a_{1},a_{2}\}$ is a basis. As in Proposition 3.5(1), we know that the projection $\pi_{X_{b}}(X_{a_{i}})$ of each axis to $X_{b}$ has length at most $|b|_{\mathcal{S}}$. We claim that, furthermore, if the projections are disjoint then the arc connecting $\pi_{X_{b}}(X_{a_{1}})$ to $\pi_{X_{b}}(X_{a_{2}})$ is length at most $|b|_{\mathcal{S}}$. This claim clearly gives a uniform bound on the difference between $[a_{1}]_{b}$ and $[a_{2}]_{b}$, and with a similar case-by-case analysis as in Proposition 3.5 one can check that $[a_{1}]_{b}$ and $[a_{2}]_{b}$ differ by at most $2$.
@@ 3.2 Change of Basis
Let $T^{\mathcal{S}}$ and $T^{\mathcal{T}}$ be the Cayley trees of $F_{N}$ given by $\mathcal{S}$ and $\mathcal{T}$, respectively. Let $f:T^{\mathcal{S}}\to T^{\mathcal{T}}$ be the map that sends each edge of $T^{\mathcal{S}}$ to the geodesic in $T^{\mathcal{T}}$ with the same endpoints. Let $1_{\mathcal{S}}$ and $1_{\mathcal{T}}$ be the basepoints of the respective Cayley trees and use $g_{\mathcal{S}}=g.1_{\mathcal{S}}$ and $g_{\mathcal{T}}=g.1_{\mathcal{T}}$ to denote the vertices in the trees given by $g\in F_{N}$. If $p=[1_{\mathcal{S}},g_{\mathcal{S}}]$ is a path in $T^{\mathcal{S}}$ from $1_{\mathcal{S}}$ to the axis of an element $a\in F_{N}$, then the bounded backtracking property (see [Coo87, GJLL98]) implies that $g_{\mathcal{T}}=f(g_{\mathcal{S}})$ lies at distance at most $C$ from the axis of $a$ in $T^{\mathcal{T}}$, where $C$ is the bounded backtracking constant. Therefore, as $|b|_{\mathcal{S}}=|b|_{\mathcal{T}}$, the proposition reduces to showing that for any path $p$ based at $1_{\mathcal{S}}$, the reduction of $f(p)$ spends approximately the same amount of time travelling along the axis of $b$ in $T^{\mathcal{T}}$ as the path $p$ does along the axis of $b$ in $T^{\mathcal{S}}$ (in the same direction). This is reasonably standard: decompose such a path $p$ as $p=p_{1}p_{2}p_{3}$, where $p_{1}=[1_{\mathcal{S}},b_{\mathcal{S}}^{k}]$ for some $k$, the path $p_{2}$ is a segment of length $\leq|b|_{\mathcal{S}}$, and $p_{3}$ is a path disjoint from the axis $X^{\mathcal{S}}_{b}\subset T^{\mathcal{S}}$ except for its initial endpoint. As $f$ is bi-Lipschitz on the vertex set, $p_{3}$ can further be decomposed as $p_{3}^{\prime}p_{3}^{\prime\prime}$, where $f(p_{3}^{\prime\prime})$ is disjoint from $X^{\mathcal{T}}_{b}$ and $p_{3}^{\prime}$ is of uniformly bounded length. Then the intersection of the reduction of $f(p)$ with $X^{\mathcal{T}}_{b}$ in $T_{\mathcal{T}}$ is $[1_{\mathcal{T}},b_{\mathcal{T}}^{k}]$, up to cancellation/concatenation with a subpath of $f(p_{2}p_{3}^{\prime})$, which is of uniformly bounded length. ∎
@@ 4 Building quasi-flats
In Sections 1-3 of [BH92], it is shown that an automorphism $\phi_{0}$ as above has a stable train track representative $f\colon G\to G$ on a graph $G$ with exactly one indivisible Nielsen path (iNP) $\rho$ which, as the automorphism is geometric, forms a loop representing $b$ based at a point $x\in G$ that is fixed by $f$. Furthermore, $\rho$ crosses every edge in $G$ exactly twice (see [BH92, Lemma 3.9] and the discussion in [BH92, Section 4]).
We identify $\pi_{1}(G,x)$ with $F_{N}$ via a choice of maximal tree $T$ in $G$ together with an orientation of the edges in $G\smallsetminus T$. Each edge-loop based at $x$ determines a word in the letters ${\mathcal{S}}^{\pm 1}$ whose length is the number of edges of $G\smallsetminus T$ crossed by the loop. In particular, $|b|_{\mathcal{S}}=2N$. Since $b$ is the boundary of a surface of Euler characteristic $1-N$, this tells us that $\mathcal{S}$ minimizes the length of $b$. We define $\phi$ to be the automorphism induced by $f_{*}\colon\pi_{1}(G,x)\to\pi_{1}(G,x)$.
We now prove the second part with $r\geq 0$. There exists a legal loop $l$ in $G$ based at $x$ that crosses some edge exactly once (this is a variant of Francaviglia and Martino’s Sausage Lemma [FM11, Lemma 3.14]). Any such loop represents a primitive element in $F_{N}$. As in Section 2 of [BH92], we lift $f$ to a map $\tilde{f}\colon\tilde{G}\to\tilde{G}$ on the universal cover that represents $\phi$ (i.e. $\tilde{f}(gy)=\phi(g)\tilde{f}(y)$ for all $y\in\tilde{G}$). The map $\tilde{f}$ fixes a lift $\tilde{x}$ of our basepoint $x\in G$, and the axis of $b$ (acting as a deck transformation of $\tilde{G}$) contains $\tilde{x}$. As the loop $l$ is based at $x$ its axis $\operatorname{Axis}_{\tilde{G}}(l)$ in $\tilde{G}$ also contains $\tilde{x}$. As $l$ is legal and $\tilde{x}$ is fixed by $\tilde{f}$, the axis of $\phi^{r}(l)$ also contains $\tilde{x}$ for all $r\geq 0$. It follows that iterated images of $l$ are cyclically reduced, so that $[\phi^{r}(l)]_{b}=0$ for all $r\geq 0$. For $k\in\mathbb{Z}$, the axis of ${\rm{ad}}_{b}^{k}(l)$ is also legal and contains the point $b^{k}\tilde{x}$, so that $k-1\leq[{\rm{ad}}_{b}^{k}(l)]_{b}\leq k+1$. As $b^{k}\tilde{x}$ is also fixed under $\tilde{f}$, this implies that $k-1\leq[\phi^{r}{\rm{ad}}_{b}^{k}(l)]_{b}\leq k+1$ for all $k\in\mathbb{Z}$ and $r\geq 0$. As the map $\mathcal{AF}_{N}\to\mathbb{Z}$ induced by $[A]\mapsto[A]_{b}$ is 2-Lipschitz, we have
$\begin{split}|[\phi^{r}{\rm{ad}}_{b}^{k}(A)]_{b}-[{\rm{ad}}_{b}^{k}(A)]_{b}|\leq{}&|[\phi^{r}{\rm{ad}}_{b}^{k}(A)]_{b}-[\phi^{r}{\rm{ad}}_{b}^{k}(l)]_{b}|+|[\phi^{r}{\rm{ad}}_{b}^{k}(l)]_{b}-[{\rm{ad}}_{b}^{k}(l)]_{b}|\\ &+|[{\rm{ad}}_{b}^{k}(l)]_{b}-[{\rm{ad}}_{b}^{k}(A)]_{b}|\end{split}$
[…]
for all $r\geq 0$ and $k\in\mathbb{Z}$.
@@
To deal with the case where $r<0$ in (2), we run the entire argument again for $[\phi_{0}]^{-1}$. This provides a representative $\psi$ of $[\phi_{0}]^{-1}$ fixing $b$, such that for all $A\in\mathcal{AF}_{N}$ there exists a constant $C_{A}^{\prime}$ such that $|[\psi^{r}{\rm{ad}}_{b}^{k}(A)]_{b}-[{\rm{ad}}_{b}^{k}(A)]_{b}|\leq C_{A}^{\prime}$ for all $k\in\mathbb{Z}$ and $r\geq 0$. Note that although you might be working with a different basis $\mathcal{S}^{\prime}$ on which $b$ is minimal length, the projections $\mathcal{AF}_{N}\to\mathbb{Z}$ are coarsely equivalent by Proposition 3.7. Furthermore, this is the unique representative of $[\phi_{0}]^{-1}$ fixing $b$ such that $b^{\infty}$ and $b^{-\infty}$ are nonattracting under $\partial\psi$. This implies that $\psi=\phi^{-1}$, so we can finish part (2) by taking $M_{A}:=\max\{C_{A},C_{A}^{\prime}\}$.
@@
Let $\Sigma$ be a compact surface with one boundary component, fix an identification $\pi_{1}\Sigma=F_{N}$, let $b\in F_{N}$ be the homotopy class of the boundary loop, let $\phi\in{\rm{Aut}}(F_{N})$ be the automorphism induced by a pseudo-Anosov element of $\text{{\rm{Mod}}}(\Sigma)$, let $\Lambda=\langle\phi,\,{\rm{ad}}_{b}\rangle<{\rm{Aut}}(F_{N})$ and note that $\Lambda\cong\mathbb{Z}^{2}$. Then, for every $A\in\mathcal{AF}_{N}$, the orbit map $\lambda\mapsto\lambda(A)$, defines a quasi-isometric embedding $\Lambda\cong\mathbb{Z}^{2}\hookrightarrow\mathcal{AF}$.
@@
Note that it is enough to prove the result for one orbit. We may also replace $\phi$ with ${\rm{ad}}_{b}^{k}\phi$ without changing $\Lambda$, so we can assume that $\phi$ and the basis ${\mathcal{S}}$ are as described in Proposition 4.1. Let $s\in\mathcal{S}$ and let $A=\langle s\rangle$. Then $k-1\leq[{\rm{ad}}_{b}^{k}(A)]_{b}^{\mathcal{S}}\leq k+1$ for all $k\in\mathbb{Z}$. We define a map $\mathcal{OF}_{N}\to\mathbb{Z}$ as follows: for each conjugacy class $[B]$ of a free factor we pick a closest point $[\phi]^{j}([A])$ in the $[\phi]$-orbit of $[A]$. This assignment $[B]\mapsto j$ is Lipschitz [BF14], and by composing it with the natural projection $\mathcal{AF}_{N}\to\mathcal{OF}_{N}$ we obtain a Lipschitz map $R\colon\mathcal{AF}_{N}\to\mathbb{Z}$ which is invariant under $\text{Inn}(F_{N})$. It follows from Theorem 3.6 that the map $[\,\cdot\,]_{b}^{\mathcal{S}}\times R\colon\mathcal{AF}_{N}\to\mathbb{Z}^{2}$ is also Lipschitz, and we claim that $B\mapsto\phi^{R(B)}{\rm{ad}}_{b}^{[B]_{b}}(A)$ gives a Lipshitz retraction of $\mathcal{AF}_{N}$ onto the orbit of $\Lambda=\langle\phi,{\rm{ad}}_{b}\rangle$. To see this, note that $R(\phi^{j}{\rm{ad}}_{b}^{k}(A))=j$ for all $j,k\in\mathbb{Z}$, and